OpenStudy (anonymous):

Laren is paid 1.5 times her hourly rate for overtime. She gets paid overtime if she works over 40 hours per week. Last week, she worked 48 hours and earned $468. Write an equation to tin laren's hourly rate.

OpenStudy (anonymous):

40r + 8(1.5)r = 468 40r + 12r = 468 52r = 468 r = 468/52

OpenStudy (anonymous):

is the awnser 9?

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Good luck in all of your studies and thx for the recognition! @ifrah34